# TenonOS: A Self-Generating Intelligent Embedded Operating System Framework for Edge Computing

### v1

@@ 8 Future Work
This work takes initial steps in building a modular micro-library system for TenonOS, but further efforts are needed to improve its scalability and hardware coverage.
First, the current set of reusable micro-libraries remains small. Increasing library coverage will enable broader functionality and greater architectural flexibility. Future work will focus on identifying and refactoring additional components into independent, reusable modules. Second, TenonOS currently targets a narrow range of hardware platforms. To support heterogeneous deployment scenarios, we plan to expand compatibility across diverse processors, accelerators, and edge devices. Addressing these areas will enhance TenonOS’s adaptability and extend its applicability to a wider range of embedded and real-time systems.

### v2

@@ 7 Future Work
This work takes initial steps in building a modular micro-library system for TenonOS, but further efforts are needed to improve its scalability and hardware coverage. First, the current set of reusable micro-libraries remains small. Increasing library coverage will enable broader functionality and greater architectural flexibility. Future work will focus on identifying and refactoring additional components into independent, reusable modules. Second, TenonOS currently targets a narrow range of hardware platforms. To support heterogeneous deployment scenarios, we plan to expand compatibility across diverse processors, accelerators, and embedded devices. Addressing these areas will enhance TenonOS’s adaptability and extend its applicability to a wider range of embedded and real-time systems.